OpenStudy (anonymous):

medal to the person who can help me!! Find cos theta if sin theta=0.56

OpenStudy (lynfran):

whats 0.56 in fraction form?

OpenStudy (anonymous):

14/25

OpenStudy (lynfran):

now do u know the Pythagorean triple (7,24,25)?

OpenStudy (anonymous):

no i don’t….

OpenStudy (anonymous):

If \[\sin (θ)=\frac{ opposite side }{ hypotenuse }=\frac{ 14 }{ 25 }\] and \[\cos(θ)=\frac{ adjacentside }{ hypotenuse }=\frac{ x }{ 25 }\] and it is a right triangle, so \[14^{2}+x ^{2}=25^{2}\] then \[x=\sqrt{429}\] so \[\cos (θ)=\frac{ \sqrt{429} }{ 25 }\]

OpenStudy (anonymous):

soo..is 25 the answer then? i thought i had to use the pythagorean identity..

OpenStudy (lynfran):

no i made an error ...mistook 14 for 24...lol

OpenStudy (anonymous):

ohhh so i had to find the triple to put into the equation.. GOT IT

OpenStudy (anonymous):

thnx xx :)

OpenStudy (lynfran):

@Natriumhydrid correct